OpenStudy (anonymous):

Which equation represents the line passing through the points (1, 1) and (−2, 13) 4x − y = 5 x − 4y = 5 x + 4y = 5 4x + y = 5

OpenStudy (phi):

one way is try (1,1) in each. replace x with 1 and y with 1 4x − y = 5 becomes 4*1- 1 =4-1=3 3 ≠ 5, so that one does not work.

OpenStudy (anonymous):

So what do I do next Phi, and thanks for helping me.

OpenStudy (anonymous):

Thanks Krishna! :)

OpenStudy (anonymous):

the best way to do this problem is find slope of line passing through (1,1) and (-2,13), then compare that with the options, i.e. compare the slope with the equations of lines given, once a match is found just plug in either (1,1) or (-2,13) to confirm the answer